OpenStudy (anonymous):

the product of two consecutive integers is 5 less than 5 times their sum. Find the integers.

OpenStudy (anonymous):

n (n + 1) + 5 = 5 (n + n + 1) n= 0 or n=9 0, 1 is a solution 0 (1) + 5 = 5(1+0) 9, 10 is a solution 9 (10)+5=95 = 5(9+10)